OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: 2x + 2y = 10 x = –y + 5 A. There are an infinite number of solutions. B. The solution is (4, –1). C. The solution is (2, –3). D. There is no solution.

OpenStudy (texaschic101):

2x + 2y = 10 --- reduce by dividing by 2 x + y = 5 x = -y + 5 same as the other equation...so infinite solutions

OpenStudy (anonymous):

Thank you so much!

OpenStudy (texaschic101):

YW so much...lol